Assuming the sun to be a spherical body of radius R at a temperature of T K, evaluate the total radiant power, incident on Earth, at a distance r from the Sun. ?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Assuming the sun to be a spherical body of radius R at a temperature of T K,the total radiant power, incident on Earth, at a distance r from the Sun is given by,
\frac{\pi ro^{{2}}}{4\pi r^{2}}(\sigma T^{4}4\pi R^{2})
\frac{\pi ro^{{2}}}{r^{2}}(\sigma T^{4}R^{2})
Thanks and Regards
Shaik Aasif
askIITians faculty